Please help with the problem
Find an exponential function of the form f(x)= ba^x + c with y-intercept 2, horizontal asymptote y = -2, that passes throgh the point P(1,4).

Answer by stanbon(48558)

Find an exponential function of the form f(x)= ba^x + c with y-intercept 2, horizontal asymptote y = -2, that passes throgh the point P(1,4).
-----------
y-intercept of 2 means it passes thru (0,2)
It also passes thru (1,4)
Find a and find b:
horizontal asymptote of y=-2 means c=-2
-----------
Using (0,2) and c=-2
2 = ba^0 -2
4=b
--------
Using (1,4) and c=-2 and b=4
4 = 4[a^1]-2
a=3/2
-----------
EQUATION:
y = 4[3/2)]^x -2
==========
Cheers,
Stan H.